The average weight of 25 chocolates bars selected from a normally distributed population is 200 g with a standard deviation of 10 g. Fight the point and the interval estimates using 95% confidence level.


1
Expert's answer
2022-02-25T05:17:30-0500

1. A point estimate is a single value estimate of a parameter. For instance, a sample mean is a point estimate of a population mean.


"\bar{x}=200\ g"

2. A confidence interval is the most common type of interval estimate.

The critical value for "\alpha = 0.05" and "df = n-1 = 24" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.063899."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(200-2.063899\times\dfrac{10}{\sqrt{25}},"

"200+2.063899\times\dfrac{10}{\sqrt{25}})"

"=(195.8722, 204.1278)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "195.8722<\mu<204.1278," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(195.8722, 204.1278)."